# LangYa: Revolutionizing Cross-Spatiotemporal Ocean Forecasting

### v2

@@ Abstract
Ocean forecasting is crucial for both scientific research and societal benefits. Currently, the most accurate forecasting systems are global ocean forecasting systems (GOFSs), which represent the ocean state variables (OSVs) as discrete grids and solve partial differential equations (PDEs) governing the transitions of oceanic state variables using numerical methods. However, GOFSs processes are computationally expensive and prone to cumulative errors. Recently, large artificial intelligence (AI)-based models significantly boosted forecasting speed and accuracy. Unfortunately, building a large AI ocean forecasting system that can be considered cross-spatiotemporal and air-sea coupled forecasts remains a significant challenge. Here, we introduce LangYa, a cross-spatiotemporal and air-sea coupled ocean forecasting system. Results demonstrate that the time embedding module in LangYa enables a single model to make forecasts with lead times ranging from 1 to 7 days. The air-sea coupled module effectively simulates air-sea interactions. The ocean self-attention module improves network stability and accelerates convergence during training, and the adaptive thermocline loss function improves the accuracy of thermocline forecasting. Compared to existing numerical and AI-based ocean forecasting systems, LangYa uses 27 years of global ocean data from the Global Ocean Reanalysis and Simulation version 12 (GLORYS12) for training and achieves more reliable deterministic forecasting results for OSVs. LangYa forecasting system provides global ocean researchers with access to a powerful software tool for accurate ocean forecasting and opens a new paradigm for ocean science.
@@
1 Main
@@
Ocean forecasting aims to forecast future ocean dynamic states and is crucial for understanding ocean activities. Numerical global ocean forecasting systems (GOFSs), such as the Mercator Ocean physical system (PSY4) [1] and the Real-Time Ocean Forecasting System (RTOFS) [2], rely on geophysical fluid dynamics and partial differential equations (PDEs) to describe future ocean conditions, and such systems remain widely used worldwide. However, improving the inference accuracy and speed of numerical forecast models is a significant challenge. First, numerical forecasting heavily depends on parameterizations and the extent of human knowledge regarding oceanic physical laws. Approximation functions can exacerbate forecast errors. In addition, even on supercomputers equipped with hundreds of computing nodes, it takes several hours to generate a single forecast. Large AI models are reshaping interactions between humans and technology through their massive learning parameters and powerful learning capabilities, e.g., ChatGPT and Sora. Such advances open the way for the development of geophysical system forecasting, including large AI models for weather (e.g., FourcastNet, PanGu-Weather) and ocean (e.g., XiHe [3], AI-GOMS [4]) forecasts. These large models are mainly data-driven, recasting air state variables (ASVs) and ocean state variables (OSVs) forecasting problems as autoregressive video generation, thereby avoiding much computational complexity and uncertainties inherent in physical processes. Moreover, specialized computing devices such as graphics processing units (GPUs) offer extremely fast inference speeds. By investigating and practicing ocean forecasting, we summarize two key barriers in current large ocean forecasting models: (1) Future OSVs are forecasted autoregressively in a step-by-step manner, with temporal information not incorporated into the training and inference processes. It neglects the time-dependent evolution of OSV fields and their interdependencies, making it impossible to achieve cross-timescale forecasts. (2) Air and ocean data are concatenated as model inputs to represent air-sea interactions, a simple way that cannot effectively simulate these interactions nor integrate their underlying mechanisms into the training and inference processes. Although researchers agree that large AI models hold tremendous potential for ocean forecasting, it should be acknowledged that revolutionary breakthroughs are still needed before an AI ocean forecasting system capable of cross-timescale, air-sea coupled forecasts can be realized. Fortunately, the breakthroughs have now come as expected. Here, we present the LangYa forecasting system (see the Methods section for "LangYa"), a powerful large AI-based model for cross-spatiotemporal, air-sea coupled ocean forecasting. LangYa delivers faster and more accurate forecasts for OSVs (temperature, salinity, and currents) than both traditional numerical weather prediction (NWP) systems and XiHe (an influential large AI-based model). Our contributions are fourfold: (1) We propose a large-language-model-based (LLM-based) Time Embedding Module that achieves high spatiotemporal resolution (1/12°, 1 day) forecasts for OSVs. It enables cross-timescale forecasts (up to 7 days) without iterative steps, allowing training and inference for future 7-day OSVs with one-off. (2) We develop an asynchronous cross-iterative random sampling strategy to simulate the atmospheric’s stochastic interaction with the ocean, improving the forecasting skill for variables with different rates of change, such as sea temperature and currents. (3) We introduce an ocean self-attention module based on cosine attention, enabling more stable model training and faster convergence. With this module, a single epoch of distributed training on global OSVs from 1993 to 2019 can be completed in just 0.5 days. (4) We propose an adaptive loss function for thermocline, which significantly improves the accuracy of forecasts at a 100 m depth thermocline, yielding about a 50% improvement in flow forecast accuracy compared to other methods. Using GLORYS12 reanalysis data, we verify that LangYa can achieve stable training and convergence, and effectively yield real-time, accurate, multivariate forecasts across spatiotemporal scales, especially for the thermocline layer.
@@
2 Cross-SpatioTemporal and Air-Sea Coupled Ocean Forecasting System
LangYa adopts the idea of data-driven that builds ocean forecasting as a video prediction task. The design contains training and testing phases by using oceanic and atmospheric reanalysis data as inputs and generating forecasts for Ocean State Variables (OSVs) — such as sea temperature, salinity, and currents — through single-pass inference over 1-7 days. The system uses 11,322 samples and mitigates overfitting risks by randomly shuffling sample order at the beginning of each training iteration. The training was conducted using a distributed data-parallel (DDP) strategy on a cluster of 16 NVIDIA A800 GPUs, completed in just 14 days. The core architecture of the LangYa system is shown in Figure 1a, which integrates multiple innovative components. The input state variables include ASVs, which are u- and v-components of wind speed, temperature, and relative humidity across 1000 to 200 hPa. Sea Surface State Variables (SSSVs): u- and v-components of 10-m wind speed, sea surface temperature, and mean sea level pressure. OSVs: Temperature, salinity, and currents spanning depths of 0 to -500 m. OSVs, ASVs, and SSSVs derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) fifth-generation reanalysis dataset (ERA5) and GLORYS12 reanalysis datasets. The data is complemented by temporal information processed via a time-embedding module based on LLM (Figure 1b). The module extracts spatial features for individual time steps and temporal features across multiple time steps, and then integrates prior knowledge from the LLM to encode forecast durations. The unique fusion of spatiotemporal data supports highly efficient cross-spatiotemporal forecasting. A key innovation of LangYa is its Asynchronous Cross-Iterative Random Sampling Strategy (Figure 1c), which adaptively links ASVs, SSSVs, and OSVs to simulate air-sea interactions. Unlike traditional methods that rely on physics-based assumptions or simple data concatenation, this strategy learns spatiotemporal relationships and effectively captures air-sea variation patterns, significantly enhancing forecast accuracy without introducing additional physical constraints. The processed input data are passed through an encoder-decoder architecture equipped with an Ocean Self-Attention Module (Figure 1d). The module, based on cosine attention, ensures stable training and accelerates convergence by reducing gradient explosions common in high-dimensional data processing. Additionally, LangYa incorporates an Adaptive Thermocline Loss Function (Figure 1e) to address the challenges of forecasting thermocline mutations. The loss function is tailored to capture the steep vertical gradients and dynamic variability of the thermocline, improving forecast accuracy for this critical oceanic layer. By integrating these innovative components, LangYa has achieved several breakthroughs: It significantly reduces temporal distribution discrepancies and simplifies the complexity of air-sea coupling. It supports efficient, accurate, and stable forecasting across spatiotemporal scales. It yields state-of-the-art performance, particularly in challenging regions like thermoclines and areas with intense air-sea interactions. LangYa’s comprehensive design represents a major leap forward in AI-driven ocean forecasting, offering unprecedented efficiency and reliability for global ocean research and operational applications.
3 Experimental Setting and Main Results
@@
The study trained the LangYa forecasting system using reanalysis data (GLORYS12 and ERA5) and evaluated it with both reanalysis (GLORYS12) and observational datasets. The observational datasets are derived from the GODAE Ocean View Inter-comparison and Validation Task Team (IV-TT) Class 4 framework, which provides observation datasets from Argo, Jason-1, Jason-2, and Envisat CLS AVISO level 3 satellite altimeters. 27 years of GLORYS12 data (1993-2019) were used for training LangYa, while data from 2020-2021 were reserved for testing. LangYa forecasts 128 variables, including temperature, salinity, and zonal and meridional ocean currents across 32 depth layers. Compared with reanalysis and observational data, LangYa achieved lower RMSEs for temperature and ocean currents (both zonal and meridional) than XiHe, XiHe-Autoregression (XiHe-AR), PSY4 [1], BLUElinK OceanMAPS (BLK) [5], Global Ice Ocean Prediction System (GIOPS) [6], and Forecast Ocean Assimilation Model (FOAM) numerical models (Figure 2). Notably, the released XiHe version consists of 20 separate models, each designed to forecast OSVs for 1-10 days in either the upper or lower ocean layers. XiHe-AR refers to an autoregressive approach that generates multiday forecasts by sequentially chaining two 1-day XiHe models, resulting in higher cumulative errors. For salinity, LangYa’s forecast accuracy was comparable to XiHe and the numerical models but outperformed XiHe-AR.
@@
In addition, LangYa achieved an inference time of 1 second for forecasts of any scale (1-7 days) on a single GPU, which is over 10,000 times faster than operational numerical models. Results from the 2020-2021 test data indicate that LangYa provided "perfect" quantitative forecasts (RMSE), particularly for challenging ocean phenomena such as thermoclines. Compared to the world’s leading OSVs forecasting systems (e.g., XiHe, PSY4, BLK, GIOPS, and FOAM numerical models), LangYa significantly reduced cumulative forecast errors and demonstrated remarkable improvements in thermocline forecasting.
Both quantitative and qualitative experimental results confirm that the AI-based LangYa forecasting system for OSVs holds tremendous potential for practical applications.
Figure 3 illustrates the RMSE of LangYa’s forecasts for temperature and salinity at 1, 4, and 7 days. Spatially, LangYa achieves RMSEs below 0.3°C for temperature forecasts across most ocean regions. Areas with higher RMSEs are concentrated in the western boundary currents (e.g., the Kuroshio Extension, Gulf Stream), eastern boundary currents (e.g., the Peru Current, California Current), and the Somali Current. These regions experience intense temperature variations, making them challenging to forecast accurately using numerical models or XiHe. Nevertheless, LangYa demonstrates significant advantages in these areas. For salinity, LangYa achieves RMSEs below 0.05 psu across most ocean regions, particularly in open oceans. Higher RMSEs are observed near the equator, western boundary currents, and the Indonesian Throughflow region, where salinity is heavily influenced by complex processes such as evaporation, precipitation, river runoff, and water mass exchange. Despite these complexities, LangYa delivers accurate salinity forecasts in these challenging regions. Furthermore, the spatial distribution of LangYa’s forecast errors for both temperature and salinity remains consistent as the forecast time increases, indicating the system’s high forecast stability.
Considering the significant differences in seawater characteristics across ocean basins, the global ocean was divided into eight basins (Northwest Pacific, Northeast Pacific, North Atlantic, North Indian Ocean, South Atlantic, South Pacific, South Indian Ocean, and Arctic Ocean) to evaluate LangYa’s forecasting performance. Figure 4 presents LangYa’s forecasting accuracy for four OSVs across these eight basins over 1-7 days. The results show substantial variations in RMSEs among the different ocean basins for forecasting ocean variables. Specifically, the Arctic Ocean exhibited the lowest RMSEs for U, V, and T, indicating that LangYa’s performance in this region is the most stable and accurate. This is due to the relatively stable hydrodynamic conditions and fewer drastic changes in the Arctic. Conversely, forecasting OSVs in the North Indian Ocean presented significant challenges, as monsoons, freshwater input, and complex oceanic dynamics influence the North Indian basin. The intricate physical processes and higher uncertainties posed difficulties for LangYa’s forecasts. In contrast, LangYa demonstrated more consistent forecasting performance in the other basins, showcasing the robustness of the LangYa forecasting system.
@@
Notably, in salinity forecasts, there were significant differences in RMSEs between the North Indian Ocean and North Atlantic (relatively high values) and the two Southern Hemisphere basins (South Pacific and South Indian Ocean, relatively low values). Within the analysis in Figure 3, it can be inferred that the higher RMSEs in the North Atlantic are concentrated in the Gulf Stream extension region. The rapid salinity changes and spatial heterogeneity in this area are driven by freshwater input from the Gulf Stream and rivers. In the North Indian Ocean, the low accuracy of salinity forecasts is primarily due to the influence of complex monsoon systems and river runoff (e.g., Ganges and Brahmaputra Rivers). Additionally, the presence of coastal upwelling and the Indonesian Throughflow, which intensify water mass mixing, further complicates the forecasts in this region. These areas, characterized by complex physical processes and unique oceanic conditions, highlight the need for further development and refinement of the forecasting model.
4 Thermocline Forecast
@@
The thermocline, a transitional layer in the ocean characterized by a sharp temperature gradient, is a critical interface for energy exchange, water mass mixing, and biogeochemical processes. Typically located at depths of 100-200 meter, the thermocline exhibits complex regional and seasonal variations due to its involvement in multiple coupled physical processes, such as ocean stratification, vertical mixing, and air-sea interactions. Accurate forecasting of the thermocline is essential for advancing the understanding of these processes and their implications for global climate systems, yet it has long been a challenge due to the steep gradients and non-linear dynamics in this layer.
LangYa uses a physics-guided adaptive loss function to address these challenges by enabling the model to focus on the thermocline’s complex temperature evolution. The advancement not only improves forecasting accuracy but also provides a deeper insight into the dynamics of thermocline formation, evolution, and interaction with surrounding layers.
Experimental results, shown in Figure 5, demonstrate that LangYa consistently outperforms numerical models and existing AI-based systems, such as XiHe, in temperature profile forecasting across all forecast horizons. The model’s superior performance is particularly evident near the thermocline, where traditional approaches struggle to capture the abrupt temperature changes. By minimizing RMSE near the thermocline, LangYa not only achieves unprecedented forecasting accuracy but also enhances the ability to simulate and analyze. In regions with intense air-sea interactions and surface mixing, LangYa’s performance remains robust, highlighting its ability to incorporate the stochastic and coupled nature of mixing processes. As depth increases and the model approaches the thermocline (identified by the pronounced rise in RMSE profiles), LangYa’s advantage becomes even more pronounced. The experimental result suggests that the adaptive loss function not only captures the vertical gradients but also enhances the model’s ability to generalize across varying oceanographic conditions.
@@
The improved understanding and forecasting of thermocline dynamics enabled by LangYa provide significant implications for both scientific research and practical applications. By integrating advanced AI methods with physical knowledge, LangYa sets a new benchmark for forecasting accuracy and paves the way for more comprehensive studies on ocean stratification, climate variability, and marine resource management.
5 Discussion
LangYa forecasting system addresses the bottlenecks in cross-temporal, air-sea coupled, and thermocline forecasting. By integrating Time Embedding, air-sea interaction simulation, and adaptive loss function for thermocline, LangYa demonstrates significant advantages in forecasting OSVs across temporal scales. Compared with other forecasting systems (e.g., GOFSs and XiHe), LangYa completes global OSVs inferencing within a short period and exhibits superior performance in accuracy, timeliness, and stability, offering feasibility for real-time ocean forecasting deployment.
Traditional ocean forecasting systems rely on iterative processes to estimate future OSVs, which are computationally expensive and prone to cumulative errors. LangYa overcomes the limitations by employing a Time Embedding module, enabling forecasts for up to 7 days without iterative steps. The innovation enhances its cross-temporal forecasting capabilities and improves forecasting flexibility and efficiency. In terms of simulating air-sea interactions, LangYa utilizes an asynchronous cross-iterative random sampling strategy, successfully overcoming the simplistic concatenation of air-sea data used in existing AI-based ocean forecasting systems. The strategy effectively simulates the complex interactions between the atmosphere and ocean, significantly improving the forecasting accuracy of OSVs, such as temperature and currents. Moreover, LangYa’s training and convergence stability have been substantially enhanced through the incorporation of an ocean self-attention module. Ocean self-attention module optimizes the training process, accelerates convergence, and ensures stable forecasting outcomes.
A notable breakthrough of LangYa is its capability to forecast thermocline. The thermocline, a region in the ocean characterized by sharp temperature changes, has long been a challenging area in ocean science. By designing a loss function guided by physical knowledge, LangYa effectively captures the evolution characteristics of the thermocline, significantly improving forecasting accuracy. Experiments reveal that LangYa achieves lower RMSE in thermocline forecasts compared to existing AI-based ocean forecasting systems and numerical models, accurately reflecting the complex spatial and temporal variations of the thermocline.

### v3

@@ Abstract
Ocean forecasting is crucial for both scientific research and societal benefits. Large artificial intelligence (AI)-based models have recently boosted forecasting efficiency and accuracy. However, it remains challenging to develop a comprehensive AI-driven ocean forecasting system capable of integrating cross-spatiotemporal and atmospheric forcing. This study introduces LangYa, a cross-spatiotemporal and atmospheric forcing ocean forecasting system featuring: (1) a large-language-model-based (LLM-based) time embedding to represent forecast timescale explicitly, (2) an asynchronous cross-iterative random sampling strategy to simulate atmospheric forcing impacts on ocean processes, (3) an ocean self-attention module to enhance network stability and accelerate training convergence, and (4) an adaptive loss function to capture ocean dynamics at ocean thermocline depth ranging from tens to 300 meters. Trained on 27 years of global ocean data from the Global Ocean Reanalysis and Simulation version 12 (GLORYS12). Compared to numerical and AI-based ocean forecasting systems, LangYa enables one model to make forecasts with lead times of 1 to 7 days and achieves cross-spatiotemporal forecasting at a high resolution (1/12°, daily) in ocean temperature, salinity, and current. Specifically, LangYa realizes approximately a 50% improvement in thermocline forecast accuracy. LangYa provides ocean researchers with a powerful ocean forecasting tool and establishes a new paradigm in ocean research.
@@
Main
Ocean forecasting aims to forecast future ocean dynamic states and is crucial for understanding ocean activities [1, 2, 3, 4]. Numerical global ocean forecasting systems (GOFSs), such as the Mercator Ocean’s global ocean analysis and forecasting system (PSY4) and the Real-Time Ocean Forecasting System (RTOFS), rely on geophysical fluid dynamics and partial differential equations (PDEs) to describe future ocean conditions, and such systems remain widely used worldwide [5, 6, 7, 8, 9, 10, 11].
@@
Large artificial intelligence (AI) models are reshaping interactions between humans and technology through their massive learning parameters and powerful learning capabilities [12, 13, 14, 15], e.g., ChatGPT and Sora. Such advances open the way for the development of geophysical system forecasting [16, 17, 18, 19], including large AI models for weather (e.g., FourcastNet, PanGu-Weather [20, 21, 22, 23, 24, 25]) and ocean (e.g., XiHe, WenHai, AI-GOMS [26, 27, 28, 29]) forecasts. These large models are mainly data-driven, recasting atmosphere state variables (ASVs) and ocean state variables (OSVs), forecasting problems as autoregressive video generation, thereby avoiding much computational complexity and uncertainties inherent in physical processes. Moreover, specialized computing devices such as graphics processing units (GPUs) offer extremely fast inference speeds [30, 31, 32]. However, by investigating and practicing ocean forecasting, we summarize three key barriers in current large ocean forecasting models: (1) Future OSVs are forecasted autoregressively in a step-by-step manner, with temporal information not incorporated into the training and inference processes [33]. It neglects the time-dependent evolution of OSV fields and their interdependencies, making it impossible to achieve cross-spatiotemporal forecasts [16]. (2) Atmosphere and ocean data are concatenated as model inputs to represent their interactions, in a simple way that cannot effectively simulate the influence of atmosphere on the ocean nor integrate their underlying mechanisms into the training and inference processes [33]. (3) Available AI-based open-source large models rely on an ensemble of multi-sub-models for forecasting rather than using a single model to achieve multivariate forecasting across multiple oceanic variables.
Here, we present the LangYa forecasting system (see the Methods section for "LangYa"), a large AI-based model for cross-spatiotemporal, atmospheric forcing ocean forecasting that achieves high spatiotemporal resolution (1/12°, 1 day) forecasts. LangYa delivers faster and more accurate forecasts for OSVs (temperature, salinity, and currents) than both traditional numerical weather prediction (NWP) systems and XiHe (an influential and the only openly available AI model). Our contributions are fourfold: (1) We propose a large-language-model-based (LLM-based) Time Embedding Module, which leverages LLMs’ prior knowledge and enables cross-spatiotemporal forecasts (up to 7 days) without iterative steps, i.e., using one model to make forecasts with lead times of 1 to 7 days. (2) We develop an asynchronous cross-iterative random sampling strategy to simulate the stochastic influence of the atmosphere on the ocean, improving the forecasting skill for variables with different rates of change, such as sea temperature and currents. (3) We introduce an ocean self-attention module based on cosine attention, ensuring more stable model training and faster convergence. With this module, a single epoch of distributed training on global OSVs from 1993 to 2019 can be completed in just 0.5 days. (4) We propose an adaptive loss function for thermocline, which significantly improves the accuracy of forecasts of ocean thermocline (tens to 300 meters), yielding about a 50% improvement in thermocline forecast accuracy compared to other methods. Using GLORYS12 reanalysis data, we verify that LangYa can achieve stable training and convergence, and effectively yield real-time, accurate, multivariate forecasts across spatiotemporal scales, especially for the thermocline layer. The ablation study for our contribution is detailed in the supplementary file. Ablation experiments confirm that each module significantly contributes to LangYa’s forecasting performance, reducing errors particularly in challenging thermocline regions and complex ocean dynamics.
@@
Cross-SpatioTemporal and Air-Sea Coupled Ocean Forecasting System
LangYa adopts a data-driven idea, considering ocean forecasting as a video prediction task conditioned on oceanic and atmospheric forcing fields. The design contains training and testing phases by using oceanic and atmospheric reanalysis data as inputs and generating forecasts for Ocean State Variables (OSVs) — such as sea temperature, salinity, and currents — through single-pass inference over 1–7 days. The system uses 10585 samples (daily samples across the years 1993-2021) and mitigates overfitting risks by randomly shuffling sample order at the beginning of each training iteration. The training was conducted using a distributed data-parallel (DDP) strategy on a cluster of 16 NVIDIA A800 GPUs, completed in just 14 days. Fig. 1 illustrates the architecture (Fig. 1A) and key modules (time-embedding module in Fig. 1B, Asynchronous Cross-Iterative Random Sampling Strategy in Fig. 1C, Ocean Self-Attention Module in Fig. 1D, Adaptive Thermocline Loss Function in Fig. 1E) of LangYa, a cross-spatiotemporal and atmospheric forcing ocean forecasting system.
The core architecture of the LangYa is shown in Fig. 1A, which integrates multiple components. The input state variables include ASVs, which are u- and v-components of wind speed, air temperature, and relative humidity across 1000 to 200 hPa. Sea Surface State Variables (SSSVs): u- and v-components of 10-m wind speed, sea surface temperature, and mean sea level pressure. OSVs: Temperature, salinity, and currents spanning depths of 0 to -500 m. OSVs, ASVs, and SSSVs derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) fifth-generation reanalysis dataset (ERA5) and GLORYS12 reanalysis datasets. The data is complemented by temporal information processed via a time-embedding module based on LLM (Fig. 1B). The module extracts spatial features for individual time steps and temporal features across multiple time steps, and then integrates prior knowledge from the LLM to encode forecast durations. The unique fusion of spatiotemporal data supports highly efficient cross-spatiotemporal forecasting.
A distinctive feature of LangYa is its Asynchronous Cross-Iterative Random Sampling Strategy (Fig. 1C), which adaptively links ASVs, SSSVs, and OSVs to simulate the effect of the atmosphere on the ocean. Unlike traditional methods that rely on physics-based assumptions or simple data concatenation, this strategy learns spatiotemporal relationships and effectively captures atmosphere-ocean variation patterns, significantly enhancing forecast accuracy without introducing additional physical constraints. The processed input data are passed through an encoder-decoder architecture equipped with an Ocean Self-Attention Module (Fig. 1D). The module, based on cosine attention, ensures stable training and accelerates convergence by reducing gradient explosions common in high-dimensional data processing. To overcome inaccurate thermocline forecasts, LangYa incorporates an Adaptive Thermocline Loss Function (Fig. 1E). The loss function is tailored to capture the steep vertical gradients and dynamic variability of the thermocline, improving forecast accuracy for this critical oceanic layer.
By integrating these components, LangYa: (1) significantly reduces temporal distribution discrepancies and simplifies the complexity of atmosphere-ocean coupling. (2) supports efficient, accurate, and stable forecasting across spatiotemporal scales. (3) yields fine performance, particularly in challenging regions like thermoclines and areas with intense atmosphere-ocean interactions. LangYa’s comprehensive design represents a major leap forward in AI-driven ocean forecasting, offering unprecedented efficiency and reliability for global ocean research and operational applications.
Experimental Setting and Main Results
@@
We trained the LangYa forecasting system using reanalysis data (GLORYS12 and ERA5 [34, 35]) and evaluated it with both reanalysis (GLORYS12) and observational datasets. The observational datasets are derived from the GODAE Ocean View Inter-comparison and Validation Task Team (IV-TT) Class 4 framework [36], which provides observation datasets from Argo, Jason-1, Jason-2, and Envisat CLS AVISO level 3 satellite altimeters. 27 years of GLORYS12 data (1993–2019) were used for training LangYa, while data from 2020–2021 were reserved for testing. LangYa forecasts 128 variables, including temperature, salinity, and zonal and meridional ocean currents across 32 depth layers.
Under the same experimental environment configuration, compared with reanalysis and observational data, LangYa achieved lower RMSEs (Please refer to the Model Training and Metrics section in the Methods for the definition) for temperature and ocean currents (both zonal and meridional) than XiHe (the only openly available AI model), XiHe-Autoregression (XiHe-AR), PSY4, BLK, GIOPS, and FOAM numerical models. Fig. 2A denotes the RMSE of the u-component of ocean currents for 1-7 day forecasts. Fig. 2B is the RMSE of the v-component of ocean currents for 1-7 day forecasts. Fig. 2C indicates RMSE of the ocean temperature for 1-7 day forecasts. Fig. 2D denotes the RMSE of the salinity for 1-7 day forecasts. Notably, the released XiHe version consists of 20 separate models, each designed to forecast OSVs for 1–10 days in either the upper or lower ocean layers. XiHe-AR refers to an autoregressive approach that generates multiday forecasts by sequentially chaining two 1-day XiHe models, resulting in higher cumulative errors. For salinity, LangYa’s forecast accuracy was comparable to XiHe and the numerical models but outperformed XiHe-AR.
In addition, LangYa achieved an inference time of 1 second for forecasts of any scale (1-7 days) on a single GPU, which is over 10,000 times faster than operational numerical models. Results from the 2020–2021 test data indicate that LangYa provided reasonable quantitative forecasts (RMSE), particularly for challenging ocean phenomena such as thermoclines. Compared to the world’s leading OSVs forecasting systems, LangYa significantly reduced cumulative forecast errors and demonstrated remarkable improvements in thermocline forecasting.
@@
Fig. 3 illustrates the RMSE of LangYa’s forecasts for temperature and salinity at 1, 4, and 7 days. Spatially, LangYa achieves RMSEs below 0.3°C for temperature forecasts across most ocean regions. Areas with higher RMSEs are concentrated in the western boundary currents (e.g., the Kuroshio Extension, Gulf Stream), eastern boundary currents (e.g., the Peru Current, California Current), and the Somali Current. These regions experience large temperature variations, making them challenging to forecast accurately using numerical models or XiHe. Nevertheless, LangYa demonstrates significant advantages in these areas. For salinity, LangYa achieves RMSEs below 0.05 psu across most ocean regions, particularly in open oceans. Higher RMSEs are observed near the equator, western boundary currents, and the Indonesian Throughflow region, where salinity is heavily influenced by complex processes such as evaporation, precipitation, river runoff, and water mass exchange. Despite these complexities, LangYa delivers accurate salinity forecasts in these challenging regions. Furthermore, the spatial distribution of LangYa’s forecast errors for both temperature and salinity remains consistent as the forecast time increases, indicating the system’s high forecast stability.
@@
Fig. 4 presents LangYa’s forecasting accuracy for four OSVs across these eight basins (Northwest Pacific, Northeast Pacific, North Atlantic, North Indian Ocean, South Atlantic, South Pacific, South Indian Ocean, and Arctic Ocean) over 1–7 days. The results show substantial variations in RMSEs among the different ocean basins for forecasting ocean variables. Specifically, the Arctic Ocean exhibited the lowest RMSEs for U, V, and T, indicating that LangYa’s performance in this region is the most stable and accurate. This is due to the relatively stable hydrodynamic conditions and fewer drastic changes in the Arctic. Conversely, forecasting OSVs in the North Indian Ocean presented significant challenges, as monsoons, freshwater input, and complex oceanic dynamics influence the North Indian basin. The intricate physical processes and higher uncertainties posed difficulties for LangYa’s forecasts. In contrast, LangYa demonstrated more consistent forecasting performance in the other basins, showcasing the robustness of the LangYa forecasting system.
@@
Notably, in salinity forecasts, there were significant differences in RMSEs between the North Indian Ocean and North Atlantic (relatively high values) and the two Southern Hemisphere basins (South Pacific and South Indian Ocean, relatively low values). It can be inferred from Fig. 3 that the higher RMSEs in the North Atlantic are concentrated in the Gulf Stream extension region. The rapid salinity changes and spatial heterogeneity in this area are driven by freshwater input from the Gulf Stream and rivers. In the North Indian Ocean, the low accuracy of salinity forecasts is primarily due to the influence of complex monsoon systems and river runoff (e.g., the Ganges and Brahmaputra Rivers). Additionally, the presence of coastal upwelling and the Indonesian Throughflow, which intensify water mass mixing, further complicates the forecasts in this region. These areas, characterized by complex physical processes and unique oceanic conditions, highlight the need for further development and refinement of the forecasting model.
Thermocline Forecast
The thermocline, a transitional layer in the ocean characterized by a sharp temperature gradient, is a critical interface for energy exchange, water mass mixing, and biogeochemical processes. Typically located at tens to 100 m depths, the thermocline exhibits complex regional and seasonal variations due to its involvement in multiple coupled physical processes, such as ocean stratification, vertical mixing, and atmosphere-ocean interactions. Accurate forecasting of the thermocline is essential for advancing the understanding of these processes and their implications for global climate systems, yet it has long been a challenge due to the steep gradients and non-linear dynamics in this layer [37].
LangYa employs a physics-guided adaptive loss function specifically designed to address these challenges by directing the model’s attention to the complex temperature dynamics of the thermocline. This approach not only enhances forecasting accuracy but also deepens our understanding of thermocline formation, evolution, and its interactions with adjacent ocean layers.
The experimental results illustrated in Fig. 5 demonstrate that LangYa consistently delivers superior performance in forecasting temperature profiles across all evaluated forecast horizons, outperforming both conventional numerical models and established open-source AI-based systems, such as XiHe. LangYa’s advantage is especially pronounced near the thermocline, a region where traditional methods often struggle to accurately represent rapid temperature transitions. By minimizing RMSE near the thermocline, LangYa achieves unprecedented forecasting accuracy and significantly improves its ability to simulate and analyze ocean dynamics. In regions with intense atmosphere-ocean interactions and surface mixing, LangYa’s performance remains robust, highlighting its ability to incorporate the stochastic and coupled nature of mixing processes. As depth increases and the model approaches the thermocline (identified by the pronounced rise in RMSE profiles), LangYa’s advantage becomes even more pronounced. The experimental result suggests that the adaptive loss function not only captures the vertical gradients but also enhances the model’s ability to generalize across varying oceanographic conditions.
@@
Discussion
The LangYa forecasting system addresses critical bottlenecks in cross-spatiotemporal and atmospheric-forced ocean forecasting, as well as thermocline forecasting. By integrating Time Embedding, asynchronous cross-iterative random sampling strategy, and adaptive loss function for thermocline, LangYa demonstrates significant advantages in forecasting OSVs across temporal scales. Compared with other forecasting systems (e.g., GOFSs and XiHe), LangYa completes global OSVs inferencing within a short period and exhibits superior performance in accuracy, timeliness, and stability, offering feasibility for real-time ocean forecasting deployment.
Traditional ocean forecasting systems rely on iterative processes to estimate future OSVs, which are computationally expensive and prone to cumulative errors. LangYa overcomes the limitations by employing a Time Embedding module, enabling forecasts for up to 7 days without iterative steps. This approach enhances its cross-temporal forecasting capabilities and improves forecasting flexibility and efficiency. In terms of the effect of the atmosphere on the ocean, LangYa utilizes an asynchronous cross-iterative random sampling strategy, successfully overcoming the simplistic concatenation of atmosphere-ocean data used in existing AI-based ocean forecasting systems. The strategy effectively simulates the complex effect of the atmosphere on the ocean, significantly improving the forecasting accuracy of OSVs, such as temperature and currents. Moreover, LangYa’s training and convergence stability have been substantially enhanced through the incorporation of an ocean self-attention module. The ocean self-attention module optimizes the training process, accelerates convergence, and ensures stable forecasting outcomes.
A significant strength of LangYa is its capability to accurately forecast thermocline dynamics. The thermocline, a region in the ocean characterized by sharp temperature changes, has long been a challenging area in ocean science. By designing a loss function guided by physical knowledge, LangYa effectively captures the evolution characteristics of the thermocline, significantly improving forecasting accuracy. Experiments reveal that LangYa achieves lower RMSE in thermocline forecasts compared to existing AI-based ocean forecasting systems and numerical models, accurately reflecting the complex spatial and temporal variations of the thermocline.